OpenStudy (anonymous):

l2x+1l=7 absolute value.

OpenStudy (saifoo.khan):

2x + 1 = 7 & 2x + 1 = -7 Solve.

jimthompson5910 (jim_thompson5910):

l2x+1l=7 2x+1=7 or 2x+1=-7 2x=7-1 or 2x=-7-1 2x=6 or 2x=-8 x=6/2 or x=-8/2 x=3 or x=-4

OpenStudy (anonymous):

x=3... right?

jimthompson5910 (jim_thompson5910):

or x = -4

jimthompson5910 (jim_thompson5910):

both solutions work

OpenStudy (anonymous):

okay... so if i was doing lx-4l=9 would x=14?

jimthompson5910 (jim_thompson5910):

no, because lx-4l=9 l14-4l=9 l10l=9 10=9...which is false

OpenStudy (anonymous):

oops typo.. i mean 13?

jimthompson5910 (jim_thompson5910):

yes lx-4l=9 l13-4l=9 l9l=9 9=9...which is true there is another solution, what is it?

OpenStudy (anonymous):

lx+4l=9 so would it be 9-4=5... so x would equal 5?

jimthompson5910 (jim_thompson5910):

More like -5 since |x-4| = 9 |-5-4| = 9 |-9| = 9 9 = 9 ... which is true

OpenStudy (anonymous):

woohoo! im getting the hang of this.